OpenStudy (anonymous):

Five standard solutions of HBr are prepared by serial dilution in which, at each stage, 10.00mL are diluted to 150.00mL. Given that the concentration of the most dilute solution is 2.37x10^-6 M, determine the concentration of the original HBr stock solution

OpenStudy (anonymous):

I got .00004 M but it doesn't seem right.

OpenStudy (anonymous):

What are your steps?

OpenStudy (anonymous):

I know I have to 2.37x10^-6 x 150 mL = M x 10 mL for about 4 of them.

OpenStudy (anonymous):

I'm not sure how to solve this.

OpenStudy (anonymous):

mk

OpenStudy (anonymous):

Thanks for looking.

OpenStudy (anonymous):

@sweetburger

OpenStudy (anonymous):

I could post how someone explained it and see if it makes sense to you. If you would like.

OpenStudy (anonymous):

If you make your example by starting let's say with 0.002 mol in 0.01 L solution (which is 10 mL) and keep on diluting it to 0.15L (which is 150 mL) and calculate the final M, you'll see that it is decreasing in molarity by a factor of 15. Which makes sense because 0.15L is 15 times greater that 0.01. If you take the molarity of the diluted solution multiplied by 15 you'll get the molarity of the previous solution. So I think to get back to you're original concentration it is final concentration multiplied by 15 raised to the 5 -1 \[final M (15)^{5 -1}\] You can verify if your answer is correct by diluting 5 timesto 150 mL (0.15L)

OpenStudy (anonymous):

I get .11998125 when I do it that way.

OpenStudy (anonymous):

So take 0.112 M calculate the moles of 0.01L and then dilute (that mole) to 0.15L and repeat for a total of 5 and see if you get 2.37x10^-6.

OpenStudy (anonymous):

Where is the .112 M coming from?

OpenStudy (anonymous):

I get .11998125 rounded

OpenStudy (anonymous):

I get that too.

OpenStudy (anonymous):

Why wouldn't it be .12 M?

OpenStudy (anonymous):

You are correct. Verify if that is the correct answer.

OpenStudy (anonymous):

2.73x10^-6 *10.00/150.00 = 1.82 x 10^-7

OpenStudy (anonymous):

1.82x10^-7 *10.00/150.00 = 1.21x10^-8

OpenStudy (anonymous):

1st concentration: 0.12 M x 0.01L = 1.2x10^-3 mole 2nd concentration: (1.2x10^-3)/0.15 = 8 x 10^-3 M 3rd concentration 8 x 10^-3 x 0.01 = 8 x 10^-5 mol (8 x 10^-5) / 0.15 = 5.3333 x 10^-4 M 4th concentration 5.3333 x 10^-4 x 0.01 = 5.3333 x 10^-6 mol (5.3333 x 10^-6) / 0.15 = 3.5555 x 10^-5 M 5th concentration 3.5555 x 10^-5 x 0.01 = 3.5555 x 10^-7 (3.5555 x 10^-7) / 0.15 = ?

OpenStudy (anonymous):

2.370 x 10^-6

OpenStudy (anonymous):

Is .12M the right amount of sig figs?

OpenStudy (anonymous):

It should probably be 0.120 M. Do you submit the answer online?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Is it the correct answer?

OpenStudy (anonymous):

It counted it wrong but it makes sense though. If you can plug it back in like that.

OpenStudy (anonymous):

Is this AP chemistry or college chemistry? Can you email your professor and show the work. I think we made a slight mistake. The question might be, a stock solution (that was already prepared) was diluted 5 times. So it would be raised to the 5th power (not 5 -1). How many tries do you have?

OpenStudy (anonymous):

That was my last try. It's college.

OpenStudy (anonymous):

Can you let me know the correct answer.

OpenStudy (anonymous):

Is it a regular chem class or honors?

OpenStudy (anonymous):

If I ever find out. It doesn't tell you the right answer afterwards for some reason.

OpenStudy (anonymous):

It's a regular chem class.

OpenStudy (anonymous):

I think you should email your professor, you may get partial credit.

OpenStudy (anonymous):

I will. Thanks

OpenStudy (thadyoung):

You can use M\(_1\)V\(_1\) = M\(_2\)V\(_2\) to solve. Essentially, you're going to be doing this a few times since this was a serial dilution that you did. Or, you can simply figure out your dilution factor(s) (DF) and go from there using DF = \(\sf \frac{C_1}{C_2}\)